You have discovered a new mutation in Drosophila that causes the flies to have leg tremors. You name the mutation shaker (sh), and determine that it is an autosomal recessive mutation. You want to determine whether the sh gene is linked to the recessive mutation causing loss of wings (apterous, ap). You cross a fly heterozygous for sh and ap with a fly homozygous for both sh and ap, and obtain the following results:

sh+ ap+ 321 flies

sh ap 318 flies

sh+ ap 88 flies

sh ap+ 92 flies

a.Are the shaker and apterous genes linked? Why or why not?

b.What is the recombination frequency between these two genes?
Answer #1

The greater number are of parental type and the lowest number are of recombinant type.

Parents: sh+ ap+ / sh ap × sh ap / sh ap

Gametes sh ap
sh+ ap+ sh+ ap+/ sh ap (parental)321
sh ap sh ap/ sh ap ( parental)318
sh+ ap sh+ ap / sh ap ( recombinant)88
sh ap+ sh ap+/ sh ap ( recombinant)92

Genetic distance between two genes is {(88+92)/(88+92+318+321)}×100= (180/819)×100=21.978022 map units.

The shaker and apterous genes are linked to each other and they are found on the same chromosome. Because the genetic distance between two genes is 21.97 map units. There is not found equal proportion of each progeny. This suggests that shaker and apterous genes are not linked